# Inference Time Alignment of Diffusion Models with Evolutionary Algorithms

### v1

@@ B.2 Extended Alignment Objective Commentary
Here, we make additional comments upon Eq. 2 as presented in § 3. Recall that Eq. 2:
@@
cast inference-time alignment as search problem where we find the search distribution $q_{\phi}(\psi)$, that maximizes the expected reward. This objective is a generalization of Eq. 1 to natural evolutionary strategies [50]. Namely, we rather than searching for a single $\psi$ that maximizes the expected reward, we assume that $\psi$ is sampled from a parameterized search distribution, $q_{\phi}$, and we maximize the expected reward under this distribution. Note: $x$ is either sampled from the solution space $\psi$ in a stochastic manner or computed deterministically. In this context, $\theta$ is the parameterization of the sampler of the diffusion model, and the sampler is modeled either as $x\sim p_{\theta}(x|\psi)$ (stochastic) or $x=f_{\theta}(\psi)$ (deterministic).
@@ B.3 Extended Characterization of Evolutionary Algorithm Components
Genetic Algorithms. Genetic algorithms maintain a population of candidate solutions, $\psi_{i}$, and iteratively improve them via selection, crossover, and mutation [8]. In the context of Eq. 2, the population is viewed as an empirical distribution. Thus the objective is to maximize the expected reward of a population of solutions. This objective is maximized iteratively and at each generation:
@@
    Evaluation: For each solution $\psi_{i}$ in the population, the fitness (reward) is evaluated — i.e., $R(x_{i})$ where $x_{i}\sim p_{\theta}(x|\psi_{i})$.
@@
    Crossover: Pairs of parents exchange noise component (or transformation parameters) to produce offspring. Correct choice of crossover can help ensure we meet (C1). For example, when using uniform crossover (swapping each coordinate independently), each child’s coordinate is drawn from one of two i.i.d. $\mathcal{N}(0,1)$ parents thus remaining with high-density shell (§ B.4).
@@
Natural Evolutionary Strategies. Natural evolutionary strategies perform black-box optimization by adapting the parameterized search distribution $q_{\phi}$ over solutions [50]. Unlike GAs, $q_{\phi}$ is a parameterized search distribution (e.g., multivariate normal) and at each iteration its parameters are updated as follows:
@@
B.4 Proof of Gaussian Marginal Preservation under Uniform Crossover
@@
In this section we show that uniform crossover maintains the solutions within the high-density Gaussian shell, which was alluded to in  § 3.3. This is relevant specifically to the EvoTorch implementation of the CoSyne algorithm, which uses a uniform distribution to decide crossover behavior.
Claim
Let $X=(X_{1},\dots,X_{d})$ and $Y=(Y_{1},\dots,Y_{d})$ be independent draws from $\mathcal{N}(0,I_{d})$. To perform coordinate-wise uniform crossover, for each $i$ we independently sample:
@@
which is the CDF of $\mathcal{N}(0,I_{d})$. Equivalently, $Z\sim\mathcal{N}(0,I_{d})$. ∎
Implication
As noted in § 3.3, uniform crossover will ensure that solutions are with the high-density shell of the gaussian thus maintaining sample quality.
Appendix C Details on Parameterization of Evolutionary Algorithms
Here, we review more details on how we initialized our search parameters. We also comment on consequences and insights related to choices of initialization.
GA Mutation Rate and Stability. When using GAs, the mutation operator introduces Gaussian noise to promote exploration. However, excessive perturbations can push solutions outside the high-density shell, leading to poor sample quality. To address this, we use small mutation step sizes, which help ensure that offspring remain within the high-density shell and preserve sample quality. Typically, our mutation size is $\sigma\approx 0.1$.
Evolutionary Strategy Initialization. Here, we briefly describe pitfalls with improper ES initialization, and how we addressed them in our work. Fig. 6 visually depicts this concept. We noted that ES are particularly sensitive to initialization — in the context of this work, this "initialization" is the initial search distribution $q_{\phi}(\psi)$.
One naive, but simple way to initialize $q_{\phi}(\psi)$ is to model it as a zero-mean ($\mu=0$), isotropic (covariance $\sigma=I$) multivariate Gaussian distribution. However, even after a few optimizations steps, the ES algorithm may update $\mu$ and $\sigma$ such that sampling from $q_{\phi}(\psi)$ no longer yields samples on the Gaussian shell. As a result, sample quality is degraded significantly. This is visually depicted in the top row of Fig. 6.
@@ Appendix C Details on Parameterization of Evolutionary Algorithms
Appendix D Extended Evalaution Results
Here, we introduce additional quantiative and qualitative results.
D.1 Significance Testing
We verify whether our results are statistically significant compared to unguided, random sampling of noise by performing Wilcoxon’s signed rank test; which assumes that distribution of pair differences is symmetric. We compare the resulting $p$ value from Wilcoxon’s against the (standard) significance level $\alpha=0.05$. Entries in  Tab. 2 and  Tab. 1 are annotated with a dagger if they are statistically insignificant. Across both tables, only 8 out of 84 of the entries related to our method (CoSyne, SNES, PGPE) were insignificant.
D.2 Effect of longer optimization timelines for gradient-based methods
@@
Figures 7, 8 and 9 depict the effect of longer optimization timelines on the best-sample, mean, and median rewards for the gradient-based method DNO, compared to gradient-free methods. Gradient-based require longer optimization timelines to reach the same reward as gradient-free methods.
Figure 7: DNO with Long Optimization Horizon: Best-sample reward per step measured (ImageReward) as measured on Open Image Preferences. Horizontal lines are the maximum reward achieved by the gradient-free methods in 50 steps.
Figure 8: DNO with Long Optimization Horizon: Mean reward per step measured (ImageReward) as measured on Open Image Preferences. Horizontal lines are the maximum (mean) reward achieved by the gradient-free methods in 50 steps.
Figure 9: DNO with Long Optimization Horizon: Median reward per step measured (ImageReward) as measured on Open Image Preferences. Horizontal lines are the maximum (median) reward achieved by the gradient-free methods in 50 steps.
@@ D.3 Population Size Ablation
Figures 10, 11 and 12 depict the effect of increasing population size on the CoSyne, SNES, and PGPE algorithms. All algorithms benefit from larger population sizes, however PGPE appears to benefit the most.
@@
Figure 10: Genetic Algorithm: CoSyNE Reward statistics for the CosSyNE algorithm across optimization step.
@@
D.4 Selection Pressure and Convergence
Fig. 13 shows the effect on increasing selection pressure (tournament size) on the median reward and standard deviation. We note that high selection pressure (larger tournament size) results in a rapid reduction in the reward standard deviation as shown in Fig. 13(b), this is in line with prior work characterizing the effect of tournament selection [30].
@@ D.4 Selection Pressure and Convergence
Figure 13: Genetic Algorithm: CoSyne Reward statistics per step, as measured on Open Image Preferences. Reward statistics are measured with increasing selection pressure, i.e., increasing tournament sizes of 2, 4, and 8. We fix the population size to 16.
@@ D.5 Qualitative Results
D.5.1 PGPE JPEG Reward-Hacking
Fig. 14 shows the proclivity of PGPE to "reward-hack" as noted in § 4.2. In many instances, PGPE was able to reduce the JPEG file size, but the resulting images were far too dissimilar or nonsensical compared with the intended image for a prompt. This behavior is reward-hacking, because the algorithm completely neglects the prompt, and instead produces nonsensical images to maximize reward.
@@ D.5.2 Stable Diffusion Qualitative Results
Here, we present sets of images from DrawBench and Open Image Preferences to illustrate qualitative results — Fig. 15, Fig. 16, Fig. 17, and Fig. 18. Specifically, we show how genetic algorithms such as CoSyne feature low sample diversity, even as optimization steps scale. This can also happen to ES methods (Fig. 18, however we noted it was less likely to occur with ES. Over longer optimization horizons, ES was able to maximize rewards better than GA while having higher diversity ( § 4.4).
Figure 15: Qualitative Sample Diversity: Randomly selected DrawBench prompts evaluated on StableDiffusion-1.5 with ImageReward. CoSyne was used to perform alignment. Across optimization steps, the diversity between samples quickly diminishes. We identify this as a consistent occurrence with genetic algorithms such as CoSyne.
Figure 16: Qualitative Sample Diversity: Randomly selected Open Image Preference prompts evaluated on StableDiffusion-1.5 with ImageReward. CoSyne was used to perform alignment. Note the low sample diversity.

### v2

@@ B.2 Extended Alignment Objective Commentary
Here, we make additional comments upon Eq.˜2 as presented in Sec.˜3. Recall that Eq.˜2:
@@
cast inference-time alignment as search problem where we find the search distribution $q_{\phi}(\psi)$, that maximizes the expected reward. This objective is a generalization of Eq.˜1 to natural evolutionary strategies [wierstra2014natural]. Namely, we rather than searching for a single $\psi$ that maximizes the expected reward, we assume that $\psi$ is sampled from a parameterized search distribution, $q_{\phi}$, and we maximize the expected reward under this distribution. Note: $x$ is either sampled from the solution space $\psi$ in a stochastic manner or computed deterministically. In this context, $\theta$ is the parameterization of the sampler of the diffusion model, and the sampler is modeled either as $x\sim p_{\theta}(x|\psi)$ (stochastic) or $x=f_{\theta}(\psi)$ (deterministic).
@@ B.3 Extended Characterization of Evolutionary Algorithm Components
Genetic Algorithms.
Genetic algorithms maintain a population of candidate solutions, $\psi_{i}$, and iteratively improve them via selection, crossover, and mutation [eiben2015introduction]. In the context of Eq.˜2, the population is viewed as an empirical distribution. Thus the objective is to maximize the expected reward of a population of solutions. This objective is maximized iteratively and at each generation:
@@
    Evaluation: For each solution $\psi_{i}$ in the population, the fitness (reward) is evaluated — i.e. $R(x_{i})$ where $x_{i}\sim p_{\theta}(x|\psi_{i})$.
@@
    Crossover: Pairs of parents exchange noise component (or transformation parameters) to produce offspring. Correct choice of crossover can help ensure we meet (C1). For example, when using uniform crossover (swapping each coordinate independently), each child’s coordinate is drawn from one of two i.i.d. $\mathcal{N}(0,1)$ parents thus remaining with high-density shell (Sec.˜B.5).
@@
Natural Evolutionary Strategies.
Natural evolutionary strategies perform black-box optimization by adapting the parameterized search distribution $q_{\phi}$ over solutions [wierstra2014natural]. Unlike GAs, $q_{\phi}$ is a parameterized search distribution (e.g. multivariate normal) and at each iteration its parameters are updated as follows:
@@
B.4 Noise Transform Search Implementation
Earlier we introduced the algorithm for direct noise search ( Algorithm˜1). We noted that the algorithm for noise noise transform search is slightly different: we include it here for clarity.
Algorithm 2 Alignment via Noise Transformation Search
1:Pretrained diffusion model $p_{\theta}$, reward $R(\cdot)$, noise $z_{T}\sim\mathcal{N}(0,I)$, iterations $T$, population size $M$
2:Search variable: $\psi=A$, with transformed noise $z_{T}^{\prime}=Q(A)\,z_{T}$
3: where $Q(A)$ denotes the orthonormal component of AA obtained via QR decomposition
4:Search distribution $q_{\phi}(\psi)$:   GA: population $\{A_{i}\}_{i=1}^{M}$ or (ES) parameterized AA
5:for $t=1,\dots,T$ do
6:  Sample transformation parameters $\{A_{i}\sim q_{\phi}\}_{i=1}^{M}$
7:  Compute transformed noises $\{z_{i}^{\prime}=Q(A_{i})\,z_{T}\}_{i=1}^{M}$
8:  Generate samples $\{x_{i}\}_{i=1}^{M}$ with $x_{i}=f_{\theta}(z_{i}^{\prime})$
9:  Compute rewards $\{r_{i}\}_{i=1}^{M}$ with $r_{i}=R(x_{i})$
10:  $\phi\leftarrow\text{EAUpdate}(\phi,\{A_{i}\}_{i=1}^{M},\{r_{i}\}_{i=1}^{M})$
11:end for
12:return final $q_{\phi}$ or best transform
NB: EAUpdate depends on the specific evolutionary algorithm used. A bias term $b$ may be included, but we set it to 0.
B.5 Proof of Gaussian Marginal Preservation under Uniform Crossover
@@
In this section we show that uniform crossover maintains the solutions within the high-density Gaussian shell, which was alluded to in  Sec.˜3.2. This is relevant to the EvoTorch implementation of the CoSyne algorithm, which uses a uniform distribution to decide crossover behavior (parameterized by a probability $p$ to perform crossover).
Background.
Samples from a high-dimensional Gaussian lie on a hypersphere, this phenomenon is known as the Gaussian Annulus Theorem. Diffusion models that use Gaussian noise are (implicitly) trained under this condition [gaussian_annulus_lecture_2017]. Thus, to ensure valid generated images, we must enforce this condition (draw samples near the surface of the sphere or shell) when we perturb noise. Uniform crossover is one such pertrubation, which we now investigate.
Lemma.
Let $X=(X_{1},\dots,X_{d})$ and $Y=(Y_{1},\dots,Y_{d})$ be independent draws from $\mathcal{N}(0,I_{d})$. To perform coordinate‐wise uniform crossover, for each $i$ we independently sample:
@@
which is the CDF of $\mathcal{N}(0,I_{d})$. Equivalently, $Z\sim\mathcal{N}(0,I_{d})$.
Implication.
As noted in Sec.˜3.2, uniform crossover will ensure that solutions are with the high-density shell of the gaussian thus maintaining sample quality.
Appendix C Details on Parameterization of Evolutionary Algorithms and Compared Methods
Here, we review more details on how we initialized our EA search parameters, and how other algorithms were configured for fair comparison.
C.1 EA Parameterization
Genetic Algorithm Mutation Rate and Stability.
When using GAs, the mutation operator introduces Gaussian noise to promote exploration. However, excessive perturbations can push solutions outside the high-density shell, leading to poor sample quality. To address this, we use small mutation step sizes, which help ensure that offspring remain within the high-density shell and preserve sample quality. Typically, our mutation size is $\sigma\approx 0.1$.
Evolutionary Strategy Initialization.
Here, we briefly describe pitfalls with improper ES initialization, and how we addressed them in our work. Fig.˜7 visually depicts this concept. We noted that ES are particularly sensitive to initialization — in the context of this work, this "initialization" is the initial search distribution $q_{\phi}(\psi)$.
One naive, but simple way to initialize $q_{\phi}(\psi)$ is to model it as a zero-mean ($\mu=0$), isotropic (covariance $\sigma=I$) multivariate Gaussian distribution. However, even after a few optimizations steps, the ES algorithm may update $\mu$ and $\sigma$ such that sampling from $q_{\phi}(\psi)$ no longer yields samples on the Gaussian shell. As a result, sample quality is degraded significantly. This is visually depicted in the top row of Fig.˜7.
@@ C.1 EA Parameterization
C.2 Configuration of Compared Methods
Hyperparameter Description and Nuance.  We list the primary hyperparameters of each method in Tab.˜6; these define the configurations reported in Tab.˜7.
Although some notions—such as “batch size” BB—appear across multiple methods, they do not necessarily correspond to the same underlying mechanism. In our method, $B$ either equals the population size ($B=P$) when evaluations are fully parallelized, or is smaller than the population ($B<P$) when we process individuals in batches (Sec.˜3.3). Other methods do not necessarily treat the batch dimension in the same manner. Instead, their batch of particles or samples is more similar to a population of samples to generate rather than the number of samples they evaluate per-step. For example, DSearch-R interprets $B$ as a number of samples to output — but it resamples and replaces samples within the batch throughout the denoising process.
Choosing Configurations.  Tab.˜7 shows configurations used in Tabs.˜1 and 2 . It was not straightforward to map our hyperparameters to those of other methods, given the differences noted above. To ensure a fair comparison, we therefore selected configurations that maximized alignment capability while matching the per-prompt running time of our strongest method, CoSyNE, on each reward function.
Table 6: Hyperparameters of Surveyed Methods. This is a companion table for Tab.˜7. We refer the reader to the papers (and respective codebases) for a more detailed explanation of each. Batch size $B$ is interpreted differently by each method, per Sec.˜C.2.
   Algorithm(s)    Parameter    Description
   CoSyNE, SNES, PGPE    $P$    Population size
   $B$    Batch size, parallel evaluations
   $O$    Optimization Steps
   FKS[singhal_fk_steering_2025]    PC    Particle Count, similar to population size $P$
   $t_{\text{start}}$    Resampling denoising step start
   $t_{\text{end}}$    Resampling denoising step end
   $t_{\text{freq}}$    Resampling frequency
   DSearch-R[li2025dynamic]    $B$    Similar to population size $P$
   $n_{\text{duplicates}}$    Beam width
   $w$    Tree expansion factor
   SVDD[li_svdd_2024]    $B$    Similar to population size $P$
   $n_{\text{duplicates}}$    Beam width
Table 7: Tabs.˜1 and 2 Hyperparameter Configurations. These configurations closely match the running time of our best method, CoSyNE for each alignment objective. HPSv2 is slower to evaluate than all other alignment objectives, partially due to its large CLIP backbone. We note that batch size $B$ is interpreted differently by each method, per Sec.˜C.2.
Algorithm(s) Objective Configuration
FKS[singhal_fk_steering_2025] ImageReward, CLIP, JPEG $\text{PC}=128$, $t_{\text{start}}=5$, $t_{\text{end}}=45$, $t_{\text{freq}}=1$
SVDD[li_svdd_2024] ImageReward, CLIP, JPEG $B=6$, $n_{\text{duplicates}}=20$
DSearch-R[li2025dynamic] ImageReward, CLIP, JPEG $B=8$, $n_{\text{duplicates}}=5$, $w=5$, $r=0.125$
FKS HPSv2 $\text{PC}=16$, $t_{\text{start}}=5$, $t_{\text{end}}=45$, $t_{\text{freq}}=1$
SVDD HPSv2 $B=6$, $n_{\text{duplicates}}=10$
DSearch-R HPSv2 $B=8$, $n_{\text{duplicates}}=4$, $w=4$, $r=0.125$
CoSyNE, SNES, PGPE ImageReward, CLIP, HPSv2, JPEG $B=16$, $P=16$
Appendix D Extended Evaluation Results
D.1 Cross-Model Evaluation on Open Image Preferences
Table 8: Open Images Preferences Cross-Model Evaluation The best‐sample reward and the median reward of the best population achieved by each algorithm on Open Image Preferences across three models: StableDiffusion 1.5, StableDiffusion 3, and PixArt-$\alpha$. Algorithms are configured in the same manner as in  Tab.˜1. Higher rewards are better.
Algorithm SD1.5 SD3 PixArt-$\alpha$ (LCM)
CLIP ImgReward CLIP ImgReward CLIP ImgReward
Best Med. Best Med. Best Med. Best Med. Best Med. Best Med.
Best-of-N $39.1$ $34.9$ $1.49$ $0.59$ $39.0$ $36.0$ $1.72$ $1.38$ $38.8$ $35.9$ $1.67$ $1.25$
Zero–Order $39.5$ $36.4$ $1.54$ $0.98$ $39.4$ $37.4$ $1.75$ $1.60$ $38.9$ $36.1$ $1.67$ $1.29$
Ours: Direct Noise Search
SNES $40.1$ $37.6$ $1.57$ $1.25$ $39.9$ $38.1$ $1.77$ $1.68$ $38.9$ $35.8$ $1.68$ $1.21$
PGPE $39.0$ $36.1$ $1.37$ $0.88$ $38.4$ $36.4$ $1.67$ $1.50$ $38.7$ $35.8$ $1.65$ $1.23$
[…]
By virtue of our results, we show that optimizing noise with EAs is often sufficient to perform alignment (see Tab.˜1), though its effectiveness varies by model (Tab.˜8). This raises two questions: (1) Are some latent spaces easier to search than others? (2) How can one design, train, or modify diffusion models to make their latent spaces easier to search? Our reward improvements on PixArt-$\alpha$ were diminished compared to other diffusion models (Tab.˜8), suggesting some property of the latent space and/or diffusion model that affects our search efficacy.
D.2 Effect of Longer Optimization Horizons on DNO
@@
Figures˜8(a), 8(b) and 8(c) depict the effect of longer optimization timelines on the best-sample, mean, and median rewards for the gradient-based method DNO, compared to gradient-free methods. Gradient-based require longer optimization timelines to reach the same reward as gradient-free methods.
@@ D.3 Population Size Ablation
Figures˜9, 10 and 11 depict the effect of increasing population size on the CoSyNE, SNES, and PGPE algorithms. All algorithms benefit from larger population sizes, however PGPE appears to benefit the most.
D.4 Selection Pressure and Convergence
Fig.˜12 shows the effect on increasing selection pressure (tournament size) on the median reward and standard deviation. We note that high selection pressure (larger tournament size) results in a rapid reduction in the reward standard deviation as shown in Fig.˜12(b), this is in line with prior work characterizing the effect of tournament selection [miller1995genetic].
(a) Best-sample reward per step. Horizontal lines show the max reward in 50 steps.
(b) Mean reward per step. Horizontal lines show the max mean in 50 steps.
(c) Median reward per step. Horizontal lines show the max median in 50 steps.
Figure 8: DNO with Long Optimization Horizon. ImageReward per step on Open Image Preferences.
@@ D.4 Selection Pressure and Convergence
Figure 9: Genetic Algorithm: CoSyNE Reward statistics for the CoSyNE algorithm across optimization step.
@@
Figure 12: Genetic Algorithm: CoSyNE Reward statistics per step, as measured on Open Image Preferences. Reward statistics are measured with increasing selection pressure, i.e. increasing tournament sizes of 2, 4, and 8. We fix the population size to 16.
@@ D.5 Qualitative Results
D.5.1 Reward Hacking in Sec.˜4.2
Here we discuss reward-hacking behavior on our method, PGPE, and minor instances across other surveyed methods. Fig.˜17 shows the proclivity of PGPE to "reward-hack" as noted in Sec.˜4.2. In many instances, PGPE was able to reduce the JPEG file size, but the resulting images were far too dissimilar or nonsensical compared with the intended image for a prompt. This behavior is reward-hacking, because the algorithm completely ignores key prompt details, and instead produces nonsensical images to maximize reward.
@@ D.5.2 Stable Diffusion Qualitative Results
Here we illustrate the differences between ES and GA in longer optimization horizons. We show reward statistics in Fig.˜13, Fig.˜14, Fig.˜15, and Fig.˜16. GAs like CoSyNE feature low sample diversity, even as optimization steps scale. This can also happen to ES methods (Fig.˜16, however we noted it was less likely to occur with ES. Over longer optimization horizons, ES was able to maximize rewards better than GA while having higher diversity (Sec.˜4.3).
Figure 13: Qualitative Sample Diversity: Randomly selected DrawBench prompts evaluated on StableDiffusion-1.5 with ImageReward. CoSyNE was used to perform alignment. Across optimization steps, the diversity between samples quickly diminishes. Genetic algorithms such as CoSyNE are particularly vulnerable to this phenomenon.
Figure 14: Qualitative Sample Diversity: Randomly selected Open Image Preference prompts evaluated on StableDiffusion-1.5 with ImageReward. CoSyNE was used to perform alignment. Note the low sample diversity across optimization steps. We identify this as a consistent occurrence with genetic algorithms such as CoSyne.
@@ D.6 Extended DrawBench Results
Here we include additional, granular results on the DrawBench dataset. We provide this data in Tab.˜10, Tab.˜11, Tab.˜12, and Tab.˜13. Specifically, we include additional statistics (min, max, median, mean, standard) for each reward function. Each subtable for a particular reward function displays the population-best rewards and median rewards, which were measured across all prompts of the DrawBench dataset.
D.7 Total Reward Function Evaluations
The cost of evaluating the reward function is significantly higher than the cost of sampling [uehara2025inference]e.g. human feedback. Therefore, we report the number of total reward function evaluations from Tabs.˜1 and 2, which we illustrate in Tab.˜9. The evaluation costs are relatively low for the alignment objectives in this work. Even so, a method that can achieve high rewards with fewer samples is desirable in any context.
We show our methods took substantially fewer reward evaluations than similar works FKD, SVDD, and DSearch-R, while achieving higher aesthetic scores than all other methods.
For clarity, we describe how we derived the evaluation counts that appear in  Tab.˜9, which are based on the configurations from Tab.˜7.
Best-of-N, Zero-Order.  These methods were parameterized by $P=16$, run over 15 optimization steps for 240 total evaluations. Each sample in the population was evaluated once per optimization step.
CoSyNE.  CoSyNE generates an intermediate population size that is $1.5\times$ its input and output population size $P$, as mentioned in Sec.˜4.5. Given $P=16$, an effective population of $P=24$ is realized; over 15 optimization steps this leads to 360 total evaluations.
FKS.  FKS resampled latents a total of 41 times (denoising step 5 until step 45) over a population of 128 particles for 5258 total evaluations. Each particle is evaluated once at each denoising step.
SVDD.  SVDD maintains 120 candidate latent noise vectors over 49 denoising steps for 5880 total evaluations. Each candidate is evaluated once at each denoising step.
DSearch-R.  DSearch-R maintains 5 candidate latent noise vectors, and follows a schedule to further expand its search, yielding $\sim$7880 evaluations under their default (exponential) search schedule. This resamples latents at $\sim$38 denoising steps.
Table 9: Number of reward function evaluations per-prompt performed by each method from Tabs.˜1 and 2. If a reward evaluation was batched ($B>1$), we count that as $B$ evaluations.
   Algorithm    Reward Evaluations
   Best-of-N    240
   Zero-Order    240
   FKS    5258
   SVDD    5880
   DSearch-R    7880
   CoSyNE    360